OpenStudy (anonymous):

simplify (2/x^2+8x+15) + (1/x^2+11x+30)

OpenStudy (anonymous):

show work please

OpenStudy (anonymous):

combine like terms a start is \[\frac{1}{x^2}+\frac{2}{x^2}=\frac{3}{x^2}\]

OpenStudy (anonymous):

\[(\frac{2}{x^2}+8x+15)+(\frac{1}{x^2}+11x+30)\] \[\frac{3}{x^2}+19x+45\] Multiply all terms by the LCD and use the LCD as denominator \[\frac{3+19x^3+45x^2}{x^2}\]